EXERCISE 2 Show that two portfolios on the capital allocation line are perfectly correlated EXERCISE 3
Answer #1

Two portfolios on the capital market line have the same slope but different standard deviation.

The Capital allocation line equation is given by

E(r_{p})=r_{f}+\left ( \frac{E(r_{a})-r_{f}}{\sigma _{a}} \right )\sigma _{p}

Here, \left ( \frac{E(r_{a})-r_{f}}{\sigma _{a}} \right ) is the slope of the CML and is same for both the portfolios

Lets us consider a case of 2-portfolio on a CML with different portfolio standard deviation \sigma _{a}

E(r_{a}) is the expected return on the risky-part of the portfolio

r_{f} is the risk-free rate = 2%

\sigma _{a} is the standard deviation on risky-part of the portfolio = 0.4

\sigma _{p1} is the standard deviation of portfolio 1 = 0.2

\sigma _{p2} is the standard deviation of portfolio 2 = 0.25

We use excel, to get different expected return on the portfolios E(r_{p1}) and E(r_{p2}) using the risk-free rate and standard deviations mentioned

E(Rp1) 0.02+0.2*((E(Ra)-0.02)/0.4)
E(Rp2) 0.02+0.25*((E(Ra)-0.02)/0.4)
E(Ra) E(Rp1) E(Rp2)
1% 1.500% 1.37500%
2% 2.000% 2.00000%
3% 2.500% 2.62500%
4% 3.000% 3.25000%
5% 3.500% 3.87500%
6% 4.000% 4.50000%
7% 4.500% 5.12500%
8% 5.000% 5.75000%
9% 5.500% 6.37500%
10% 6.000% 7.00000%
11% 6.500% 7.62500%
12% 7.000% 8.25000%
13% 7.500% 8.87500%
14% 8.000% 9.50000%
15% 8.500% 10.12500%
16% 9.000% 10.75000%
17% 9.500% 11.37500%
18% 10.000% 12.00000%
19% 10.500% 12.62500%
20% 11.000% 13.25000%
21% 11.500% 13.87500%
22% 12.000% 14.50000%
23% 12.500% 15.12500%
24% 13.000% 15.75000%
25% 13.500% 16.37500%
26% 14.000% 17.00000%
27% 14.500% 17.62500%
28% 15.000% 18.25000%
29% 15.500% 18.87500%
30% 16.000% 19.50000%
31% 16.500% 20.12500%
32% 17.000% 20.75000%
33% 17.500% 21.37500%
34% 18.000% 22.00000%
35% 18.500% 22.62500%
36% 19.000% 23.25000%
37% 19.500% 23.87500%
38% 20.000% 24.50000%
39% 20.500% 25.12500%
40% 21.000% 25.75000%
41% 21.500% 26.37500%
42% 22.000% 27.00000%
43% 22.500% 27.62500%
44% 23.000% 28.25000%
45% 23.500% 28.87500%

Now, we calculated the correlation for both these portfolios using excel => Data => Data Analysis => Correlation

E F K B E(Ra) ? X Correlation Input Input Range: ок | Cancel Grouped By: C E(Rp1) 1.500% 2.000% 2.500% 3.000% 3.500% 4.000% 4

We get the correlation matrix as

E(Rp1) E(Rp2)
E(Rp1) 1
E(Rp2) 1 1

Since, the correlation between the portfolios is 1, we observe that the two portfolios (on the CML) are perfectly correlated.